OpenStudy (anonymous):

A sample of air is slowly passed through aqueous Sodium hydroxide and then over heater copper. Which gases are removed by this process? a) Carbon dioxide and water vapour b) Carbon dioxide and oxygen c) Nitrogen and oxygen d) Nitrogen and water vapour

OpenStudy (anonymous):

Question number2 The equation explains the colour change that occurs when aqeous potassium dichromate(6), what happens to the oxidation state of the chromium and the pH of the reaction mixture? Oxidation State of chromium pH of the mixture a) decrease decrease b) decrease increase c) stays the same decreases d) stays the same increases

OpenStudy (anonymous):

the equation \[K _{2}Cr_{2}O_{7} + 2KOH \rightarrow 2K_{2}CrO_{4} + H_{2}O\] potassium potassium chromate(6) dichromate(6) yellow orange

OpenStudy (callisto):

For first question. slowly passed through aqueous Sodium hydroxide => it should be acidic gas => CO2 then over heater copper. => copper + air = copper oxide => O2 So the answer is B

OpenStudy (callisto):

ON of Cr2O7^2- = +6 ON of CrO4 ^2- = +6 => ON = unchanged Since OH^-1 ions are consumed in the reaction, the solution becomes less alkaline => pH decreases So the answer is C

OpenStudy (anonymous):

Bro this was a MCQ question for my semester exam, and i wrote the same answers as u did, bt ended up with bth wrng...

OpenStudy (anonymous):

K2Cr2O7 (oxidising agent) oxidises itself and causes the reduction of other substance ( oxidises= oxidation number increase) by the definition cant we say oxidation state increases?

OpenStudy (callisto):

But there is no redox reaction there.... The ON doesn't change!!!

OpenStudy (anonymous):

but the colour changes, does it mean that it has been oxidised?

OpenStudy (callisto):

Nope, it's simply because the substances present are different!!!

OpenStudy (anonymous):

so why did the colour change?

OpenStudy (anonymous):

and why did i get it wrong..i put C as my answer and got it wrng

OpenStudy (callisto):

dichromate ion is orange in colour chromate ion is yellow in colour you can say there is a chemical conversion. But it's not a redox!!!

OpenStudy (callisto):

Is that all given in your question ?

OpenStudy (anonymous):

yeah can u explain this? ince OH^-1 ions are consumed in the reaction, the solution becomes less alkaline => pH decreases

OpenStudy (callisto):

pH value is depends on the concentration of the H+ and OH- ion As OH- is used in the reaction, its concentration decreases. Hence, the solution becomes less alkaline

OpenStudy (anonymous):

why is it used and even f it is used (KOH is added in excess) it cn replace the lost IONS and why doesnt the pH increase

OpenStudy (callisto):

you haven't mentioned in the question that it is in excess nor KOH would be added continuously...

OpenStudy (anonymous):

Okai bro...now i get it...so why s OH- used? and how?

OpenStudy (anonymous):

bro this is a part of the question? as a result of adding an excess of aqueous potassium hydroxide to aqueous potassium dichromate (6), what hapens to the oxidation state of the chromium and the pH of the reaction mixture

OpenStudy (callisto):

in the equation , KOH is one of the reactants, if the reaction takes place, of course it would be consumed :S

OpenStudy (anonymous):

hm ok but KOH is added in excess...thts in the question

OpenStudy (anonymous):

and my answer was D actually and got it wrng

OpenStudy (callisto):

even so, the pH would not increase lol, it would only decrease because OH is consumed BTW, my answer is C, not D , judge my answer again, I still can't believe my answer is wrong :P

OpenStudy (anonymous):

hehe, okai then thats the answer fine ( i just mixed this with neutralisation reaction ) so thts why i got it as pH increased. anyway thnks bro...u r a genius ur answers wnt be wrng :P

OpenStudy (callisto):

I'm not a genius and I always get wrong in chem :S

OpenStudy (anonymous):

bro i have another question can u just tell me how to work this out?

OpenStudy (anonymous):

Silver ions react with chloride ions \[Ag ^{+}(aq) + Cl^{-}(aq)\rightarrow AgCl (s)\] It is found that 5cm^3 of a 0.1 mol/dm^3 solution of the chloride metal x needs 10 cm^3 of 0.1 mol/dm^3 of silver nitrate for the complete reaction what is the formula of the chloride? \[XCl_{4}\] \[XCl_{2}\] \[XCl\] \[X_{2}Cl\]

OpenStudy (callisto):

no. of mole of metal chloride = 5/100 *0.1= 0.005 mole No. of mole of silver nitrate= 10/100 * 0.1 = 0.01 mole From the equation , no. of Cl ion used = 0.01mol Therefore there are 0.01/0.005 = 2 Cl- ions in each metal chloride the formula = XCl2

OpenStudy (anonymous):

From the equation , no. of Cl ion used = 0.01mol ?

OpenStudy (callisto):

Ag + Cl -> AgCl no of mole of Cl = no of mole of Ag = 0.01 (used in the reaction)

OpenStudy (callisto):

Ag+(aq) + Cl-(aq) -> AgCl(s)

OpenStudy (anonymous):

why dont we take metal chloride? 0.005

OpenStudy (callisto):

i just concentrate on the Cl- in that case No. of mole of Cl- used / no. of mole of metal chloride used = no of Cl- in each metal chloride

OpenStudy (anonymous):

if u dont mind can u explain it a bit further, cuz i did it in another way and gt the answer

OpenStudy (callisto):

Like MgCl2 + 2AgNO3 -> Mg(NO3)2 + 2AgCl no of mole of Ag = 0.01 no of mole of MgCl2 = 1/2 *no of mole of Ag = 0.005 Hmmm.. it's really difficult for me to explain :S

OpenStudy (anonymous):

okai its alright..thnks i thnk i get it...are u a guy or a girl?

OpenStudy (callisto):

girl !

OpenStudy (anonymous):

LOL, i was calling u a "bro" from the beginning...dint u hve the mind to corrct me?

OpenStudy (callisto):

ppl said they use 'man' to call sb... I guess it's similar ... I've thought of correcting you.. but ..anyway...

OpenStudy (anonymous):

so what are you doing (exams and stuff)?

OpenStudy (callisto):

Hmm???

OpenStudy (anonymous):

u r doing any exams? which grde are u in?

OpenStudy (callisto):

Yup exams, i assume it's grade 12 as you would call, the grade before uni.

OpenStudy (anonymous):

oh alright...is it igcse?

OpenStudy (callisto):

Nope, it's called DSE here

OpenStudy (anonymous):

so after exams engineering eh?

OpenStudy (callisto):

engineering is tough... I don't know :S

OpenStudy (anonymous):

hm nothing is easy, if u thnk it as hard u will never do it:)

OpenStudy (callisto):

Hmm.. engineering doesn't seem to be a good subject for ordinary girl like me :P

OpenStudy (anonymous):

haha...so what r u planning to do then? i guess u have the potential to work for it..cz u seem to clear my doubts which i thnk is sumthng great :) my advice is tht u go for it..

OpenStudy (callisto):

Chemistry.... It's not my cup of tea :P No exact planning at the moment. It's something after the exam :) Though, I've ranked my uni choices

OpenStudy (anonymous):

oh...something great to hear....LOL..btw u all r awarded with scholarships? or do u all pay for the whole course?

OpenStudy (anonymous):

@CAllisto u there??????????????????

OpenStudy (callisto):

Don't understand... Sorry :S